OpenStudy (anonymous):

The volume of a hexagonal pyramid is represented by the expression 12x3 cubic inches and its height is 24 inches. The area of the hexagonal base is ____x3 square inches.?

OpenStudy (anonymous):

volume of pyramid = 1/3 * height * area of base

OpenStudy (anonymous):

so plug in values of the height and volume and solve for the area

OpenStudy (anonymous):

ans. pls i cant figure out with this calc.

OpenStudy (anonymous):

ok 12x^3 = 1/3 * 24 * area = 8*area area = 12x^3 / 8 = 3x^3 / 2 or 1.5 x^3